60 –kg man lifts a 50-kg sack of rice and carries it on is shoulder along a 200-m straight path.

Find:

a) the work done by the man on the load as he lifted it from the ground to his shoulders if the distance from the ground to the shoulders is 4.0 ft.

b) the power he expended in lifting the load to his shoulders if he did this in 2.0 seconds

c) the work done on the load along the straight path

d) the power used to carry the load along the straight line if he carried it for 3.0 minutes


1
Expert's answer
2021-02-25T11:28:26-0500

(a)

"W=mgh=50\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot4\ ft\cdot\dfrac{0.3048\ m}{1\ ft}=597\ J."

(b)

"P=\dfrac{W}{t}=\dfrac{597\ J}{2\ s}=298.5\ W."

(c)

"W=Fd=mgd=50\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot200\ m=98000\ J."

(d)

"P=\dfrac{W}{t}=\dfrac{98000\ J}{3\cdot60\ s}=544.4\ W."